# Schrödinger operators with concentric $\delta$–shell interactions

### v1

@@ Abstract
We study Schrödinger operators on $\mathbb{R}^{3}$ with finitely many concentric spherical $\delta$–shell interactions. The operators are defined by the quadratic form method and are described by continuity across each shell together with the usual jump condition for the radial derivative. Using a boundary integral approach based on the free Green kernel and single–layer potentials, we derive an explicit resolvent representation for an arbitrary number of shells with bounded coupling strengths. This yields a concrete Kreĭn–type formula and a boundary operator whose noninvertibility characterizes the discrete spectrum, and it is compatible with a partial–wave reduction under rotational symmetry.
We then specialize to the two–shell case with constant couplings and obtain a detailed description of the negative spectrum. In particular, we show that the ground state (when it exists) lies in the ss–wave sector and derive an explicit secular equation for bound states. For large shell separation, each bound level approaches the corresponding single–shell level with exponentially small corrections, while a genuine tunneling splitting appears when the single–shell levels are tuned to coincide. As a simple calibration, we relate the two–shell parameters to representative core–shell quantum dot scales. At the level of order–of–magnitude and qualitative trends, Type I configurations yield a relatively strongly confined state, whereas Type II configurations produce a comparatively shallow outer–shell state.
@@ 1 Introduction
Motivated by these works, we consider Schrödinger operators on $\mathbb{R}^{3}$ with finitely many concentric spherical $\delta$–shell interactions. Let $0<R_{1}<R_{2}<\cdots<R_{N}$ and set $S_{j}:=\{x\in\mathbb{R}^{3}:\ |x|=R_{j}\}$ for $j=1,2,\ldots,N$. For each $j$ we allow a bounded measurable surface strength $\alpha_{j}\in L^{\infty}(S_{j};\mathbb{R})$. Formally, the model is
@@
where $\alpha_{j}$ acts by multiplication on the trace of a function on $S_{j}$. We define $H_{N}$ rigorously by the quadratic form method. At an abstract level, we note that resolvent representations for self–adjoint extensions are well-known. For a wide class of singular perturbations, we can express $(H_{N}-z)^{-1}$ in terms of the free resolvent and an operator–valued boundary term of Kreĭn type. For example, we refer to the extension theoretic framework in [1] and to related formulations for singular interactions in [2]. See also [23] for Kreĭn–type resolvent formulas in a general singular perturbation framework. For $\delta$– and $\delta^{\prime}$–interactions supported on hypersurfaces, including analogues of the Birman–Schwinger principle and a variant of Kreĭn’s formula, we refer to [24]. The purpose of the present paper is to derive, for concentric spherical $\delta$–shells, a concrete boundary integral version of such a formula directly in terms of the free Green kernel and single–layer potentials, without invoking boundary triple machinery, and to connect this representation to an explicit partial–wave reduction and a detailed two–shell spectral analysis. Our first main result is an explicit resolvent formula for $H_{N}$ in terms of layer potentials on the shells. This representation yields a boundary operator $K_{N}(z)$ on $\bigoplus_{j=1}^{N}L^{2}(\mathbb{S}^{2})$ such that, for $z\in\mathbb{C}\setminus[0,\infty)$, the points $z$ in the discrete spectrum correspond to the failure of invertibility of $K_{N}(z)$.
After establishing the NN–shell resolvent framework, we specialize to the two–shell case $N=2$,
@@
For the rotationally symmetric two–shell setting with constant surface strengths, one can distinguish the Type I and Type II sign patterns familiar from core–shell quantum dots. In this idealized case, Type I systems correspond to $\alpha_{1}<0<\alpha_{2}$ [11], while Type II systems correspond to $\alpha_{1}>0>\alpha_{2}$ [12]. Throughout this paper, the Type I/II terminology is used only in the constant case of $\alpha_{j}$.
@@ 4 Variational preliminaries and higher partial waves
Let $\mathcal{X}$ be a Hilbert space and let $T_{0}$ and $T$ be self–adjoint operators in $\mathcal{X}$ which are bounded from below and have the same closed form domain $Q$. Assume that the spectrum of each operator below its essential spectrum is discrete so that $\{\lambda_{k}(\cdot)\}_{k\geq 1}$ is well-defined. Denote their quadratic forms by $q_{0}$ and $q$. Assume that
@@
$\lambda_{k}(T)\geq\lambda_{k}(T_{0})\qquad\text{for all }k\geq 1,$
@@
$\lambda_{k}(T)=\sup_{\begin{subarray}{c}L\subset Q\\ \dim L=k-1\end{subarray}}\ \inf_{\begin{subarray}{c}f\in Q\\ f\perp L,\,\|f\|=1\end{subarray}}q[f],\qquad\lambda_{k}(T_{0})=\sup_{\begin{subarray}{c}L\subset Q\\ \dim L=k-1\end{subarray}}\ \inf_{\begin{subarray}{c}f\in Q\\ f\perp L,\,\|f\|=1\end{subarray}}q_{0}[f].$
Since $q\geq q_{0}$ on $Q$, the inner infima satisfy
@@
for every subspace $L\subset Q$ with $\dim L=k-1$. Taking the supremum over all such $L$ gives $\lambda_{k}(T)\geq\lambda_{k}(T_{0})$.
@@
We recall the standard partial–wave decomposition (see, for example, [1, Sec. II.2]). Let $\{Y_{\ell m}\}_{\ell\geq 0,\,-\ell\leq m\leq\ell}$ be an orthonormal basis of spherical harmonics in $L^{2}(S^{2})$. Then $L^{2}(\mathbb{R}^{3})$ decomposes as the orthogonal sum

### v2

@@ Abstract
We study Schrödinger operators on $\mathbb{R}^{3}$ with finitely many concentric spherical $\delta$–shell interactions. The operators are defined via quadratic forms and characterized by continuity across each shell together with the standard jump condition for the normal derivative. Using a boundary integral approach based on the free Green kernel and single–layer potentials, we derive an explicit resolvent formula for an arbitrary number of shells with bounded coupling strengths. This yields a concrete Kreĭn–type representation and a boundary operator whose noninvertibility characterizes the discrete spectrum, and it is compatible with partial–wave reduction under rotational symmetry.
We then specialize to the two–shell case with constant couplings and obtain a detailed description of the negative spectrum. In particular, we prove that the ground state, when it exists, lies in the ss–wave sector and derive an explicit secular equation for bound states. For large shell separation, each bound level approaches the corresponding single–shell level with exponentially small corrections, while a genuine tunneling splitting appears when the single–shell levels are tuned to coincide. As a simple calibration, we relate the two–shell parameters to representative core–shell quantum dot scales and identify the qualitative distinction between Type I and Type II configurations.
@@ 1 Introduction
The rigorous treatment of $\delta$–interactions supported on a sphere was initiated by Antoine, Gesztesy, and Shabani [4], where the single–shell case was analyzed within an extension–theoretic framework. In the specific setting of finitely many concentric spherical shells with $\delta$–type interactions, Shabani [24] investigated the model by an explicit reduction to radial one–dimensional problems. In particular, for constant shell strengths the matching conditions at the interfaces lead to a finite–dimensional determinant condition in each partial wave, so that the bound state problem can be treated in a fully explicit manner. These works provide closely related precedents for the present study, as they address essentially the same geometric configuration and exhibit, already at the level of partial–wave reduction, the characteristic coupling mechanism induced by multiple shells. While Shabani’s analysis proceeds via a reduction to one–dimensional radial ODEs and finite–dimensional matching conditions in each partial wave (in particular for constant shell strengths), our approach keeps the problem in three dimensions and yields a boundary–integral resolvent representation that remains valid for arbitrary $N$ and nonconstant surface strengths $\alpha_{j}\in L^{\infty}(S_{j};\mathbb{R})$.
Our approach is complementary. Rather than revisiting the operator–theoretic construction, we allow nonconstant strengths $\alpha_{j}\in L^{\infty}(S_{j};\mathbb{R})$ for arbitrary $N$ and derive a concrete boundary integral Kreĭn–type resolvent formula directly in terms of the free Green kernel and single–layer potentials. In the constant two–shell case we then provide a detailed large–separation analysis, including the tuned tunneling splitting on the scale $e^{-\kappa_{0}d}$. This formulation makes explicit the connection between the boundary integral picture and the partial–wave secular equations under rotational symmetry. Related developments on the scattering side for finitely many concentric sphere interactions were later obtained by Hounkonnou, Hounkpe, and Shabani [15].
Building on these works, we consider Schrödinger operators on $\mathbb{R}^{3}$ with finitely many concentric spherical $\delta$–shell interactions. Let $0<R_{1}<R_{2}<\cdots<R_{N}$ and set $S_{j}:=\{x\in\mathbb{R}^{3}:\ |x|=R_{j}\}$ for $j=1,2,\ldots,N$. For each $j$ we allow a bounded measurable surface strength $\alpha_{j}\in L^{\infty}(S_{j};\mathbb{R})$. Formally, the model is
@@
where $\alpha_{j}$ acts by multiplication on the trace of a function on $S_{j}$. We define $H_{N}$ rigorously by the quadratic form method. At an abstract level, we note that resolvent representations for self–adjoint extensions are well-known. For a wide class of singular perturbations, we can express $(H_{N}-z)^{-1}$ in terms of the free resolvent and an operator–valued boundary term of Kreĭn type. For example, we refer to the extension theoretic framework in [2] and to related formulations for singular interactions in [7]. See also [22] for Kreĭn–type resolvent formulas in a general singular perturbation framework. For $\delta$– and $\delta^{\prime}$–interactions supported on hypersurfaces, including analogues of the Birman–Schwinger principle and a variant of Kreĭn’s formula, we refer to [5]. The main result of this paper is an explicit boundary integral representation for concentric spherical $\delta$–shell interactions, written directly in terms of the free Green kernel and single–layer potentials, without invoking boundary triple machinery.
More precisely, we derive an explicit resolvent formula for $H_{N}$ in terms of layer potentials on the shells. Here $S^{2}=\{\omega\in\mathbb{R}^{3}:\ |\omega|=1\}$ denotes the unit sphere, and this representation yields a boundary operator $K_{N}(z)$ on $\bigoplus_{j=1}^{N}L^{2}(S^{2})$ such that, for $z\in\mathbb{C}\setminus[0,\infty)$, the eigenvalues of $H_{N}$ in $\mathbb{C}\setminus[0,\infty)$ are characterized by the noninvertibility of $K_{N}(z)$.
We then relate this representation to the partial–wave reduction and provide a detailed spectral analysis in the two–shell case. After establishing the NN–shell resolvent framework, we specialize to the two–shell case $N=2$,
@@
For the rotationally symmetric two–shell setting with constant surface strengths, one can distinguish the Type I and Type II sign patterns familiar from core–shell quantum dots. In this idealized case, Type I systems correspond to $\alpha_{1}<0<\alpha_{2}$ [14], while Type II systems correspond to $\alpha_{1}>0>\alpha_{2}$ [21]. This correspondence is meant only in the constant–strength setting and at the level of sign patterns. Throughout this paper, the Type I/II terminology is used only in the constant case of $\alpha_{j}$.
@@ 4 Variational preliminaries and higher partial waves
Let $\mathcal{X}$ be a Hilbert space and let $T_{0}$ and $T$ be self–adjoint operators in $\mathcal{X}$ which are bounded from below and have the same closed form domain $Q$. Assume that the spectrum of each operator below its essential spectrum is discrete so that $\{\lambda_{n}(\cdot)\}_{n\geq 1}$ is well-defined. Denote their quadratic forms by $q_{0}$ and $q$. Assume that
@@
$\lambda_{n}(T)\geq\lambda_{n}(T_{0})\qquad\text{for all }n\geq 1,$
@@
$\lambda_{n}(T)=\sup_{\begin{subarray}{c}L\subset Q\\ \dim L=n-1\end{subarray}}\ \inf_{\begin{subarray}{c}f\in Q\\ f\perp L,\,\|f\|=1\end{subarray}}q[f],\qquad\lambda_{n}(T_{0})=\sup_{\begin{subarray}{c}L\subset Q\\ \dim L=n-1\end{subarray}}\ \inf_{\begin{subarray}{c}f\in Q\\ f\perp L,\,\|f\|=1\end{subarray}}q_{0}[f].$
Since $q\geq q_{0}$ on $Q$, the inner infimum in the min–max formula satisfies
@@
for every subspace $L\subset Q$ with $\dim L=n-1$. Taking the supremum over all such $L$ gives $\lambda_{n}(T)\geq\lambda_{n}(T_{0})$.
@@
We recall the standard partial–wave decomposition (see, for example, [2, Sec. II.2]). Let $\{Y_{\ell m}\}_{\ell\geq 0,\,-\ell\leq m\leq\ell}$ be an orthonormal basis of spherical harmonics in $L^{2}(S^{2})$. Then $L^{2}(\mathbb{R}^{3})$ decomposes as the orthogonal sum
@@ 7 Tunneling Splitting of Eigenvalues
Acknowledgements
The author is grateful to Professor Pavel Exner for drawing his attention to earlier results on concentric $\delta$–shell interactions obtained by J. Shabani and collaborators. This comment led to a substantial clarification of the historical context and to an improvement of the introduction.